Find the sum of all numbers in the following array. 35+40+45+...+3000
Answer by ewatrrr(24785)

Arithmetic: d = 5
an = 35 + (n-1)5
an = 5n + 30 = 3000
n = 594
......
S(594) = (594/2)(35+3000) = 901,395